OpenStudy (anonymous):

B) 4x + 6y = -12 2x + 3y = 9 D) x - 4y = -8 2x - 3y = -16 E) x - y = -2 3x - y = 0

OpenStudy (anonymous):

Which system has (1, 3) as the solution?

OpenStudy (anonymous):

A) x + y = -1 4x + 4y = -4 this goes with it too haha

OpenStudy (anonymous):

@welshfella i thought it clicked but it didnt

OpenStudy (ashes.boom):

you can put the coordinate into each system to see which one wokrs or you can solve for x and use the solution to solve for y

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

i tried them all but it wouldnt work for me

OpenStudy (ashes.boom):

okay let me try real fast one sec

OpenStudy (ashes.boom):

okay starting with the first equation from a you subtract the y to get x by itself. x+y=-1 -y -y x=-y-1

OpenStudy (ashes.boom):

then you can put x into the next equaiton 4x + 4y = -4 4(-y-1)+4y = -4

OpenStudy (ashes.boom):

get rid of the parentheses -4y -4 + 4y = -4 -4y + 4y - 4 = -4 -> the 4y cancels out -4 = -4

OpenStudy (anonymous):

what the heck are you doing haha

OpenStudy (ashes.boom):

ugh it didnt work. basically solving the solution

OpenStudy (ashes.boom):

imma try b hold on

OpenStudy (anonymous):

C is the answer. but i used it for #1 also

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

C) x + y = -4 -3x + 2y = 2

DivineSolar (divinesolar):

Let us work out E okay @kylefoster58 ?

DivineSolar (divinesolar):

x - y = -2 3x - y = 0

OpenStudy (anonymous):

i already got it now.

DivineSolar (divinesolar):

Okay then close this question please.

OpenStudy (anonymous):

was actually coming back to do it. just seen that you said something and didnt wanna leave u hangin thank you very much!

DivineSolar (divinesolar):

Sure thing.